In how many ways can a chairperson and a vice chairperson be selected from a committee of 32 senators?
Answer by stanbon(75887)

In how many ways can a chairperson and a vice chairperson be selected from a committee of 32 senators?
------
Order is implied:
32P2 = 32!/(32-2)! = 32*31 = 992
====================================
Cheers,
Stan H.